This is copied directly from the book.
An isosceles triangle has two sides of equal length. The length of one of the equal sides is two more than 3 times the length of the third side. If the perimeter is 46m, find the length of each side.

Answer by checkley77(12844)

(3x+2)+(3x+2)+x=46
7x+4=46
7x=42
7x=42/7
x=6 ans. for the third side
3*6+2=18+2=20 for the 2 equal sides.
Proof:
20+20+6=46
46=46